Write a balanced chemical equation for each of the two different reactions. Keep in mind that the hydrogen in the carboxylic acid functional group is the acidic hydrogen in acetic acid.

0.50 M NaOH, 0.50 M HCl, phenolphthalein indicator

0.50 M NaOH, 0.50 M CH3CO2H, methyl orange indicator
Answer #1

NaOH (aq) + Hel (aq) = Naal (ar) + Hold) Naou (na) + Het (as) = Nnet (any + 140cm NaOH (aq) + CH₃COOH (aq) = CH₃COONa (ar) +